# Parallel Digital Twin-driven Deep Reinforcement Learning for User Association and Load Balancing in Dynamic Wireless Networks

### v1

@@ Abstract
Optimization of user association in a densely deployed heterogeneous cellular network is usually challenging and even more complicated due to the dynamic nature of user mobility and fluctuation in user counts. While deep reinforcement learning (DRL) emerges as a promising solution, its application in practice is hindered by high trial-and-error costs in real world and unsatisfactory physical network performance during training. In addition, existing DRL-based user association methods are usually only applicable to scenarios with a fixed number of users due to convergence and compatibility challenges. In this paper, we propose a parallel digital twin (DT)-driven DRL method for user association and load balancing in networks with both dynamic user counts, distribution, and mobility patterns. Our method employs a distributed DRL strategy to handle varying user numbers and exploits a refined neural network structure for faster convergence. To address these DRL training-related challenges, we devise a high-fidelity DT construction technique, featuring a zero-shot generative user mobility model, named Map2Traj, based on a diffusion model. Map2Traj estimates user trajectory patterns and spatial distributions solely from street maps. Armed with this DT environment, DRL agents are enabled to be trained without the need for interactions with the physical network. To enhance the generalization ability of DRL models for dynamic scenarios, a parallel DT framework is further established to alleviate strong correlation and non-stationarity in single-environment training and improve the training efficiency. Numerical results show that the proposed parallel DT-driven DRL method achieves closely comparable performance to real environment training, and even outperforms those trained in a single real-world environment with nearly 20% gain in terms of cell-edge user performance.
@@
Digital twin (DT), deep reinforcement learning (DRL), user association, load balancing, diffusion model, trajectory generation.
@@ I Introduction
In order to faithfully replicate the real-world wireless network, a DT of a wireless network should integrate not only a precise wireless channel model within the area of interest, but also an accurate user mobility model due to its significant influence on the distribution and transition pattern of channel states and base station (BS) loads. While wireless channel modeling has been extensively studied and can be implemented deterministically through measurement and ray-tracing approaches [8], the modeling of user mobility still lacks effective, efficient, and privacy-preserving methodologies [9].
Specifically, most existing network optimization works [3, 4, 10] employed random mobility models, typically the random waypoint model [11] and the Gauss Markov model [12], to represent user mobility patterns. While these models can partially simulate random user movements, their direct adherence to specific distributions can lead to a significant mismatch with real-world trajectory patterns and spatial distribution. This mismatch usually results in considerable performance degradation of DRL models in practice. Although this issue can be alleviated to some extent by employing trace-based mobility models, real user trajectories are often inaccessible due to data acquisition costs and privacy concerns [9].
With the rapid development of generative artificial intelligence (AI), various cutting-edge AI models, such as generative adversarial network (GAN) [13] and diffusion model [14], have been applied to trajectory generation tasks and achieved promising outcomes [15, 16, 17, 18, 19]. However, all these methods necessitate a significant amount of real trajectories to learn an area-specific trajectory distribution, leading to a paradox. That is, AI models struggle to generate lifelike and useful trajectories without ample real trajectories. Conversely, when ample real data is available, AI models tend to be unnecessary, as trace-based mobility models can be effectively constructed. Consequently, existing trajectory generation methods are limited to function as data augmentation and privacy protection tools for trace-based mobility models, rather than serving as genuine generative user mobility models.
To address all these issues, in this paper, we propose a high-fidelity DT construction technique with a novel generative user mobility model, and introduce a parallel DT-driven DRL method for user association and load balancing in dynamic wireless networks. The contributions of this work are summarized as follows.
@@
  • We propose a high-fidelity DT construction technique for physical wireless networks, featuring a zero-shot generative user mobility model, named Map2Traj. This model is based on diffusion models and generates lifelike trajectories solely from street maps. To the best of our knowledge, this is the first work to achieve zero-shot trajectory generation, especially for mobile terminals in wireless networks.
@@
  • The efficacy of the proposed method is validated through real-world vehicle trajectory datasets and comprehensive experiments. The evaluation involves assessing the fidelity of the constructed DT, its reliability in DRL training, and the performance of the parallel DT-driven DRL method for user association and load balancing.
@@ II-A DRL-based User Association and Load Balancing
Traditional strategies that maximize SINR in user association result in unbalanced load distribution and consequently hinder the overall performance of wireless networks. DRL was first introduced in [4] to solve the non-convex user association problems, maximizing long-term overall network utility through multi-agent DRL. Different from [4] which uses global network information, The authors in [20] trained independent agents on local observation, achieving scalable and flexible user association in a millimeter-wave network. Taking user mobility and handovers into consideration, studies in [3] and [21] employed a multi-armed bandit and a Q-learning framework, respectively, with the purposes of reducing handovers and improving throughput.
@@ II-B DT-driven Network Optimization
Despite the diversity in tasks and scenarios, these studies share a common requirement, that is, the modeling of user mobility to create DTs of wireless networks. In [22], a simple static user location model was assumed, neglecting their mobility. The random waypoint (RWP) model, which involves moving to a randomly selected destination and then choosing the next arbitrarily, was adopted by [23] and [26]. On the other hand, studies in [24] and [25] employed the Gauss Markov model (GM), characterized by using a stochastic process to model changes in user velocity and direction. Both studies in [27] and [28] employed the real-world T-Drive vehicle data set [29] to establish the aforementioned trace-based mobility model.
@@
II-C Learning-based Trajectory Generation
Initially, the learning-based trajectory generation methods were developed for synthesizing mobility data to safeguard the privacy of data providers. Liu et al. [30] first discussed the possibility of using GANs for trajectory generation, albeit without providing a detailed approach. In particular, TrajGAIL [15] employed generative adversarial imitation learning (GAIL), combining DRL and GAN, to generate trajectories through a series of next-location predictions. In [16], another approach named TrajGen was developed by transforming trajectories into images and using a deep convolutional GAN (DCGAN) to generate virtual trajectory images. In [17], TS-TrajGen was proposed to integrate GAN with the mobility analysis method, including the A* algorithm and mobility yaw reward, to enhance the model performance. DiffTraj in [18] applied a diffusion model to generate synthetic trajectories while preserving spatial-temporal features extracted from real trajectories. These studies have demonstrated commendable performance in generating privacy-preserving synthetic trajectories.
However, these methods fall short when it comes to zero-shot trajectory generation for unobserved new areas. Specifically, TrajGAIL [15] simply sampled actions from generated action probability distribution and constructs trajectory autoregressively, without the capacity to introduce data from new areas. Although TrajGen [16] used street map data to filter and calibrate generated trajectories through map matching [31], the generated trajectories adhered to the training set distribution, rather than that in new areas. Building further upon TrajGen, TS-TrajGen [17] utilized street maps to select and construct the best continuous trajectory with the A* algorithm. This process, however, remained confined to trajectories that adhere to the original distribution. Alternatively in a conditional generation manner, DiffTraj [18] employed the diffusion model and incorporated prior knowledge of trip data, such as the travel time, average speed, and distance. While these complementary knowledge do improve generation performance, they do not enable the transfer of trajectory generation to new areas.
@@ III System Model
$r_{ij}[n]=\frac{c_{ij}[n]}{\ell_{j}[n]}\left(1-\frac{T_{\mathrm{HO},i}[n]}{T_{s}}\right),$ (3)
@@
$x_{ij}[n]=\begin{cases}1,&\text{ user }i\text{ associates to BS }j\text{ in time slot }n,\\ 0,&\text{ otherwise. }\end{cases}$ (4)
@@
$\sum_{j\in\mathcal{B}}x_{ij}[n]=1,\quad\forall i\in\mathcal{U}.$ (5)
The logarithmic utility function $\log(r_{ij}[n])$ is adopted as the optimization target to balance the sum-rate and fairness, which naturally achieves load-balancing [34]. Now, the optimization task, i.e., the network utility maximization problem can be formulated as
@@ IV Distributed DRL Method for User Association
$R_{i}[n]=$ $\alpha\sum_{j\in\mathcal{B}}x_{ij}[n]\log\left(r_{ij}[n]\right)+$
$\frac{1-\alpha}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}\sum_{k\in\mathcal{U}}x_{kj}[n]\log\left(r_{kj}[n]\right),\ \alpha\in\left[0,1\right].$ (8)
The $\alpha$ is a hyper-parameter to balance between the convergence rate and overall performance. Notice that $r_{ij}[n]$ in (8) is derived from $\operatorname{SINR}_{ij}[n]$ (part of $\boldsymbol{\Phi}_{i}[n]$ from $\mathbf{s}_{i}[n]$) and $\ell_{j}[n]$ (part of $\boldsymbol{\ell}[n]$ from $\mathbf{s}_{i}[n+1]$). In addition, $x_{ij}[n]$ in (8) is an element of the action vector $\mathbf{a}_{i}[n]$. Therefore, given the transmit power $P_{j}$ and bandwidth $W_{j}$ for each $j\in\mathcal{B}$, as well as handover interruption time $T_{\mathrm{HO},i}$, the reward for the $i$-th DRL agent can be written in a function as
@@
$\underset{\pi_{\boldsymbol{\theta}}}{\operatorname{max}}\underset{\begin{subarray}{c}i\in\mathcal{U}\\ \mathbf{s}_{i}[n]\sim\mathcal{D}\\ \mathbf{a}_{i}[n]\sim\pi_{\boldsymbol{\theta}}(\cdot|\mathbf{s}_{i}[n])\\ \mathbf{s}_{i}[n+1]\sim\mathbb{P}(\cdot|\mathbf{s}_{i}[n],\mathbf{a}_{i}[n])\end{subarray}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}R_{i}[n]\right],$ (10)
@@
Figure 2: Actor-critic DRL model structure.
Since all DRL agents share the same parameters and weights, training samples $(\mathbf{s}_{i}[n],\ \mathbf{a}_{i}[n],\ \mathbf{s}_{i}[n+1],\ R_{i}[n]),\ \forall i\in\mathcal{U}$ are not distinguished by subscripts and are uniformly used for training. The training strategy here follows the state-of-the-art proximal policy optimization (PPO) algorithm [35], with several modifications to the structure of the actor network. As illustrated in Fig. 2, we integrate a long short-term memory (LSTM) layer in the neural network to capture the dynamic of states and incorporate memory capability. In addition, the input $\boldsymbol{\Phi}_{i}[n]$ is utilized to construct an action mask that selects top-$N$ BSs with the strongest signal, thereby assisting the convergence.
@@ V Parallel DT-driven DRL Method
V-A Construction of DT
To address the high trial-and-error costs, poor network performance before convergence, and limitations of the single training environment, it is crucial to develop a faithful DT environment for effective DRL training. By aligning the number of BSs and users, the state space $\mathcal{S}$ and action space $\mathcal{A}$ can be directly mirrored from the real environment. The training process for DRL agents in such a DT environment can be expressed in a way similar to (10) as
$\underset{\pi_{\boldsymbol{\theta}}}{\operatorname{max}}\underset{\begin{subarray}{c}i\in\mathcal{U}\\ \mathbf{s}_{i}[n]\sim\mathcal{D^{\prime}}\\ \mathbf{a}_{i}[n]\sim\pi_{\boldsymbol{\theta}}(\cdot|\mathbf{s}_{i}[n])\\ \mathbf{s}_{i}[n+1]\sim\mathbb{P^{\prime}}(\cdot|\mathbf{s}_{i}[n],\mathbf{a}_{i}[n])\end{subarray}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}R_{i}[n]\right],$ (11)
@@ V-A Construction of DT
While channel modeling has been extensively studied and can be implemented deterministically through measurement and ray-tracing methods [8], there are thus far few methods capable of efficiently modeling user mobility within a newly given area [9]. The random mobility models can hardly capture the distribution and trajectory patterns of real-world users due to the adherence to the stochastic process. Trace-based methods and existing learning-based trajectory generation approaches ask for a large amount of data, constrained by the high data acquisition costs and privacy issues. Inspired by the zero-shot image generation [36], which enables the creation of images from descriptions unseen during training, we devise a zero-shot trajectory generation approach to construct the generative user mobility model.
The key to zero-shot learning is to associate observed and unobserved objects through some form of auxiliary information, which encodes the inherent properties of objects [37]. The question arises: Can we apply zero-shot generation techniques to trajectory generation? More specifically, is there some form of data that is both readily accessible and closely related to real trajectories, akin to the relationship between texts and images? The answer lies in street maps, which are usually open-source [^1] https://www.openstreetmap.org/ and exhibit strong correlations with user trajectories and distribution, especially the high dynamic vehicle users, as depicted in Fig. 3. Consequently, in this study, the objects in zero-shot learning are trajectories following different area-specific distributions, while the auxiliary information is the street map. By understanding the relationship between trajectories and maps through extensive training data, the proposed Map2Traj model generates lifelike trajectories for unobserved areas piloted by their street maps, that is, zero-shot trajectory generation.
@@
Now we give a formal definition of this zero-shot trajectory generation task conditioned on street maps: Given a training set includes a set of $X$ street maps =$\mathcal{M}=\{\boldsymbol{m}^{1},\boldsymbol{m}^{2},\dots,\boldsymbol{m}^{X}\}$ and the corresponding set of real-world trajectories $\mathcal{T}=\{\mathcal{L}^{1},\mathcal{L}^{2},\dots,\mathcal{L}^{X}\}$, where each $\mathcal{L}^{x}=\{\boldsymbol{l}^{x,1},\boldsymbol{l}^{x,2},\dots,\boldsymbol{l}^{x,Y}\}$ is a set of $Y$ real trajectories within the area of street map $\boldsymbol{m}^{x}$. The objective of zero-shot trajectory generation is to develop a generative model trained on this training set $\langle\mathcal{M},\mathcal{T}\rangle$. With an unobserved street map $\boldsymbol{m}^{o}\notin\mathcal{M}$, this developed model should be capable of generating synthetic trajectories such that:
  • Each generated trajectory closely resembles some real trajectory $\boldsymbol{l}\in\mathcal{L}^{o}$ in terms of trajectory pattern, where $\mathcal{L}^{o}=\{\boldsymbol{l}^{o,1},\boldsymbol{l}^{o,2},\dots,\boldsymbol{l}^{o,Y}\}$ is a set of real trajectories within this unobserved area.
  • The set of generative trajectories matches well with real trajectories set $\mathcal{L}^{o}$ in terms of spatial distribution.
@@
In this study, trajectories $\boldsymbol{l}$ and street maps $\boldsymbol{m}$ are both processed into 192$\times$192 binary images. Although a single binary image can effectively convey the spatial layout of a street map, it falls short in depicting distinct characteristics of various road types. In the OpenStreetMap dataset, roads are tagged with attributes such as Motorway, Primary, and Residential. To exploit these attributes, we segregate roads into multiple groups, create binary images for each, and merge them into a multi-channel binary image. Due to the computational complexity of the diffusion model, we simplify this process by dividing the roads into two groups: major and minor roads.
As a tractable solution for the zero-shot trajectory generation task, our proposed Map2Traj approach exploits the diffusion model [14], consisting of a forward diffusion process and a reverse diffusion process (denoising).

### v2

# Large Vision Model-Enhanced Digital Twin with Deep Reinforcement Learning for User Association and Load Balancing in Dynamic Wireless Networks

@@ Abstract
Optimization of user association in a densely deployed cellular network is usually challenging and even more complicated due to the dynamic nature of user mobility and fluctuation in user counts. While deep reinforcement learning (DRL) emerges as a promising solution, its application in practice is hindered by high trial-and-error costs in real world and unsatisfactory physical network performance during training. Also, existing DRL-based user association methods are typically applicable to scenarios with a fixed number of users due to convergence and compatibility challenges. To address these limitations, we introduce a large vision model (LVM)-enhanced digital twin (DT) for wireless networks and propose a parallel DT-driven DRL method for user association and load balancing in networks with dynamic user counts, distribution, and mobility patterns. To construct this LVM-enhanced DT for DRL training, we develop a zero-shot generative user mobility model, named Map2Traj, based on the diffusion model. Map2Traj estimates user trajectory patterns and spatial distributions solely from street maps. DRL models undergo training in the DT environment, avoiding direct interactions with physical networks. To enhance the generalization ability of DRL models for dynamic scenarios, a parallel DT framework is further established to alleviate strong correlation and non-stationarity in single-environment training and improve training efficiency. Numerical results show that the developed LVM-enhanced DT achieves closely comparable training efficacy to the real environment, and the proposed parallel DT framework even outperforms the single real-world environment in DRL training with nearly 20% gain in terms of cell-edge user performance.
@@
Digital twin (DT), large vision model (LVM), deep reinforcement learning (DRL), user association, load balancing, diffusion model, trajectory generation.
@@ I Introduction
In order to faithfully replicate the real-world wireless network, a DT of a wireless network should integrate not only a precise wireless channel model within the area of interest, commonly referred to as the radio map [8], but also an accurate user mobility model due to its significant influence on the distribution and transition pattern of channel states and base station (BS) loads. While the modeling of wireless channel has been extensively studied and can be implemented deterministically through measurement and ray-tracing approaches [9], the modeling of user mobility still lacks effective, efficient, and privacy-preserving methodologies [10]. Specifically, most existing network optimization works [3, 4] employed random mobility models, typically the random waypoint model [11] and the Gauss Markov model [12], to represent user mobility patterns. While these models can partially simulate random user movements, their adherence to specific distributions can lead to a significant mismatch with real-world trajectory patterns and spatial distribution. This mismatch usually results in considerable performance degradation of DRL models in practice. Although this issue can be alleviated to some extent by employing trace-based mobility models, real user trajectories are often inaccessible due to data acquisition costs and privacy concerns [10].
The rapid advancement of artificial intelligence (AI) has enabled transformative applications across diverse domains, driven by breakthroughs in large-AI models. Large language models (LLMs) like GPT-4 [13] and DeepSeek [14] excel in text-based tasks, while large vision models (LVMs) such as stable diffusion [15] and Sora [16] demonstrate remarkable capabilities in visual content generation. Recently, LVMs, particularly generative adversarial networks (GANs) [17] and diffusion models [18], have been applied to trajectory generation tasks and achieved promising outcomes [19, 20, 21, 22, 23]. However, existing methods necessitate a significant amount of real trajectories to learn a trajectory distribution in a given area, leading to a paradox. That is, AI models struggle to generate lifelike and useful trajectories without ample real trajectories. Conversely, when ample real data is available, AI models tend to be unnecessary, as trace-based mobility models can be readily constructed. Therefore, existing trajectory generation methods are limited to functioning as data augmentation and privacy protection tools for trace-based mobility models, rather than serving as genuine generative user mobility models.
To address all these issues, in this paper, we introduce an LVM-enhanced DT for wireless networks featuring a novel zero-shot generative user mobility model, along with a parallel DT-driven DRL framework for user association and load balancing in dynamic wireless networks. The contributions of this work are summarized as follows.
@@
  • We propose an LVM-enhanced DT construction technique for physical wireless networks, featuring a zero-shot generative user mobility model, named Map2Traj. This user mobility model is based on LVM, specifically the diffusion model, and generates lifelike trajectories solely from street maps. To our knowledge, this is the first work to achieve zero-shot trajectory generation, especially for mobile terminals in wireless networks.
@@
  • The efficacy of the proposed method is validated through real-world vehicle trajectory datasets and comprehensive experiments. The evaluation involves assessing the fidelity of the LVM-enhanced DT, its reliability in DRL training, and the performance of the parallel DT-driven DRL method for user association and load balancing.
@@ II-A DRL-based User Association and Load Balancing
Traditional strategies that maximize SINR in user association result in unbalanced load distribution and consequently hinder the overall performance of wireless networks. DRL was first introduced in [4] to solve the non-convex user association problems, maximizing long-term overall network utility through multi-agent DRL. Different from [4] that uses global network information, the authors in [24] trained independent agents on local observation, achieving scalable and flexible user association in a millimeter-wave network. Taking user mobility and handovers into consideration, studies in [3] and [25] employed a multi-armed bandit and a Q-learning framework, respectively, with the purpose of reducing handovers and improving throughput.
@@ II-B DT-driven Network Optimization
Despite the diversity in tasks and scenarios, these studies share a common requirement, that is, the modeling of user mobility to create DTs of wireless networks. In [26], a simple static user location model was assumed, neglecting their mobility. The random waypoint (RWP) model, which involves moving to a randomly selected destination and then choosing the next one arbitrarily, was adopted by [27] and [30]. On the other hand, studies in [28] and [29] employed the Gauss Markov model (GM), characterized by using a stochastic process to model changes in user velocity and direction. Both studies in [31] and [32] employed the real-world T-Drive vehicle data set [33] to establish the aforementioned trace-based mobility model.
@@
II-C LVM-based Trajectory Generation
Initially, generative AI and LVMs were introduced to trajectory generation primarily for synthesizing mobility data, thereby safeguarding the privacy of data providers. Liu et al. [34] first discussed the possibility of using GANs for trajectory generation, albeit without providing a detailed approach. After that, TrajGAIL [19] was developed by employing generative adversarial imitation learning (GAIL), which combines both DRL and GAN, to generate trajectories through a series of next-location predictions. In [20], another approach named TrajGen was developed by transforming trajectories into images and using a deep convolutional GAN (DCGAN) to generate virtual trajectory images. In [21], TS-TrajGen was proposed to integrate GAN with the mobility analysis method, including the A* algorithm and mobility yaw reward, to enhance the model performance. DiffTraj in [22] applied a diffusion model to generate synthetic trajectories while preserving spatial-temporal features extracted from real trajectories. These studies have demonstrated commendable performance in generating privacy-preserving synthetic trajectories.
However, these methods fall short when it comes to zero-shot trajectory generation for unobserved new areas. Specifically, TrajGAIL [19] simply samples actions from generated action probability distributions and constructs trajectories autoregressively, without the capacity to introduce data from new areas. Although TrajGen [20] used street map data to filter and calibrate generated trajectories through map matching [35], the generated trajectories adhered to the training set distribution, rather than that in new areas. Building further upon TrajGen, TS-TrajGen [21] utilized street maps to select and construct the best continuous trajectory with the A* algorithm. This process, however, remained confined to trajectories that adhere to the original distribution. Alternatively, in a conditional generation manner, DiffTraj [22] employed the diffusion model and incorporated prior knowledge of trip data, such as the travel time, average speed, and distance. While this complementary knowledge does improve generation performance, it does not enable the transfer of trajectory generation to new areas.
@@ III System Model
$r_{ij}[n]=(c_{ij}[n]/\ell_{j}[n])(1-T_{\mathrm{HO},i}[n]/T_{s})$ (3)
@@
$x_{ij}[n]=\begin{cases}1,&\text{ user }i\text{ associates to BS }j\text{ in time slot }n,\\ 0,&\text{ otherwise. }\end{cases}$
@@
$\sum_{j\in\mathcal{B}}\nolimits x_{ij}[n]=1,\quad\forall i\in\mathcal{U}.$ (4)
The logarithmic utility function $\log(r_{ij}[n])$ is adopted as the optimization target to balance the sum-rate and fairness, which naturally achieves load-balancing [38]. Now, the optimization task, i.e., the network utility maximization problem, can be formulated as
@@ IV Distributed DRL Method for User Association
$R_{i}[n]=\alpha\sum\nolimits_{j\in\mathcal{B}}x_{ij}[n]\log\left(r_{ij}[n]\right)$
$+$ $\frac{1-\alpha}{|\mathcal{B}|}\sum\nolimits_{j\in\mathcal{B}}\sum\nolimits_{k\in\mathcal{U}}x_{kj}[n]\log\left(r_{kj}[n]\right),\ \alpha\in\left[0,1\right].$ (7)
The $\alpha$ is a hyperparameter to balance between the convergence rate and overall performance. Notice that $r_{ij}[n]$ in (7) is derived from $\operatorname{SINR}_{ij}[n]$ (part of $\boldsymbol{\Phi}_{i}[n]$ from $\mathbf{s}_{i}[n]$) and $\ell_{j}[n]$ (part of $\boldsymbol{\ell}[n]$ from $\mathbf{s}_{i}[n+1]$). In addition, $x_{ij}[n]$ in (7) is an element of the action vector $\mathbf{a}_{i}[n]$. Therefore, given the transmit power $P_{j}$ and bandwidth $W_{j}$ for each $j\in\mathcal{B}$, as well as handover interruption time $T_{\mathrm{HO},i}$, the reward for the $i$-th DRL agent can be written in a function as
@@
$\underset{\pi_{\boldsymbol{\theta}}}{\operatorname{max}}\underset{\begin{subarray}{c}i\in\mathcal{U},\ \mathbf{s}_{i}[n]\sim\mathcal{D}\\ \mathbf{a}_{i}[n]\sim\pi_{\boldsymbol{\theta}}(\cdot|\mathbf{s}_{i}[n])\\ \mathbf{s}_{i}[n+1]\sim\mathbb{P}(\cdot|\mathbf{s}_{i}[n],\mathbf{a}_{i}[n])\end{subarray}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}R_{i}[n]\right],$ (9)
@@
Since all DRL agents share the same parameters and weights, training samples $(\mathbf{s}_{i}[n],\ \mathbf{a}_{i}[n],\ \mathbf{s}_{i}[n+1],\ R_{i}[n]),\ \forall i\in\mathcal{U}$ are not distinguished by subscripts and are uniformly used for training. The training strategy here follows the state-of-the-art proximal policy optimization (PPO) algorithm [39], with several modifications to the structure of the actor network. Specifically, we integrate a long short-term memory (LSTM) layer in the neural network to capture the dynamics of states and incorporate memory capability. In addition, the input $\boldsymbol{\Phi}_{i}[n]$ is utilized to construct an action mask that selects top-$N$ BSs with the strongest signal, thereby assisting the convergence.
@@ V Parallel DT-driven DRL Method
V-A Construction of the LVM-enhanced DT
To address the high trial-and-error costs, poor network performance before convergence, and limitations of the single training environment in traditional DRL methods, it is crucial to develop a faithful DT environment for effective DRL training. By aligning the number of BSs and users, the state space $\mathcal{S}$ and action space $\mathcal{A}$ can be mirrored from the real environment. The training process for DRL agents in such a DT environment can be expressed in a way similar to (9) as
$\underset{\pi_{\boldsymbol{\theta}}}{\operatorname{max}}\underset{\begin{subarray}{c}i\in\mathcal{U},\ \mathbf{s}_{i}[n]\sim\mathcal{D^{\prime}}\\ \mathbf{a}_{i}[n]\sim\pi_{\boldsymbol{\theta}}(\cdot|\mathbf{s}_{i}[n])\\ \mathbf{s}_{i}[n+1]\sim\mathbb{P^{\prime}}(\cdot|\mathbf{s}_{i}[n],\mathbf{a}_{i}[n])\end{subarray}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}R_{i}[n]\right],$ (10)
@@ V-A Construction of the LVM-enhanced DT
While channel modeling has been extensively studied and can be implemented deterministically through measurement and ray-tracing methods [9], there are, thus far, few methods capable of efficiently modeling user mobility within a newly given area [10]. The random mobility models can hardly capture the distribution and trajectory patterns of real-world users due to their adherence to the stochastic process. Trace-based methods and existing learning-based trajectory generation approaches ask for a large amount of data, constrained by the high data acquisition costs and privacy issues. Inspired by the zero-shot image generation [40], which enables the creation of images from descriptions unseen during training, we devise a zero-shot trajectory generation approach to construct the generative user mobility model.
The key to zero-shot learning is to associate observed and unobserved objects through some form of auxiliary information, which encodes the inherent properties of objects [41]. The question arises: Can we apply zero-shot generation techniques to trajectory generation? More specifically, is there some form of data that is both readily accessible and closely related to real trajectories, akin to the relationship between texts and images? The answer lies in street maps, which are usually open-source and exhibit strong correlations with user trajectories and distribution, especially the high dynamic vehicle users, as depicted in Fig. 2. Consequently, in this study, the objects in zero-shot learning are trajectories following different area-specific distributions, while the auxiliary information is the street map. By understanding the relationship between trajectories and maps through extensive training data, the proposed Map2Traj model generates lifelike trajectories for unobserved areas piloted by their street maps, that is, zero-shot trajectory generation.
@@
Now we give a formal definition of this zero-shot trajectory generation task conditioned on street maps: Given a training set includes a set street maps =$\mathcal{M}=\{\boldsymbol{m}^{1},\boldsymbol{m}^{2},\dots\}$ and the corresponding set of real-world trajectories $\mathcal{T}=\{\mathcal{L}^{1},\mathcal{L}^{2},\dots\}$, where each $\mathcal{L}^{x}=\{\boldsymbol{l}^{x,1},\boldsymbol{l}^{x,2},\dots\}$ is a set of real trajectories within the area of street map $\boldsymbol{m}^{x}$. The objective of zero-shot trajectory generation is to develop a generative model trained on this training set $\langle\mathcal{M},\mathcal{T}\rangle$. With an unobserved street map $\boldsymbol{m}^{o}\notin\mathcal{M}$, this developed model should be capable of generating synthetic trajectories such that:
  • Each generated trajectory closely resembles some real trajectory $\boldsymbol{l}\in\mathcal{L}^{o}$ in terms of trajectory pattern, where $\mathcal{L}^{o}=\{\boldsymbol{l}^{o,1},\boldsymbol{l}^{o,2},\dots\}$ is a set of real trajectories within this unobserved area.
  • The set of generative trajectories matches well with the real trajectory set $\mathcal{L}^{o}$ in terms of spatial distribution.
@@
In this study, trajectories $\boldsymbol{l}$ and street maps $\boldsymbol{m}$ are both processed into 192$\times$192 binary images. Although a single binary image can effectively convey the spatial layout of a street map, it falls short in depicting distinct characteristics of various road types. In the OpenStreetMap dataset, roads are tagged with attributes such as Trunk, Primary, and Residential. To exploit these attributes, we categorize roads into multiple groups, create binary images for each group, and merge them into a multi-channel binary image.
As a tractable solution for the zero-shot trajectory generation task, our proposed Map2Traj approach exploits the LVM, specifically the diffusion model [18], to generate user trajectories. It consists of a forward diffusion process and a reverse diffusion process (denoising).